OpenStudy (anonymous):

In a problem it says that x is "at most -5." What does this mean?

OpenStudy (anonymous):

@pooja195

OpenStudy (kidrah69):

http://prntscr.com/6n8qk8 my account is broke but im pooja LOL

OpenStudy (kidrah69):

see if the screenshot helps :)

OpenStudy (anonymous):

nope wnt let me says i cant access it. im on a school computer and its blocked

OpenStudy (kidrah69):

its a screen shot at most means this symbol \[\le \]

OpenStudy (anonymous):

u can use <= or => as greater than or equal to or less than or equal but i need help

OpenStudy (kidrah69):

\[\le \] =at most \[\ge \]=atleast

OpenStudy (anonymous):

thanks how ur account mssed up

OpenStudy (kidrah69):

its stuck in a loop script dont worry about it im sure they will have it fixed soon :) and you're welcome